OpenStudy (anonymous):

derivative of y=(-4x)/y , can anyone help me?

OpenStudy (anonymous):

\[\frac{ -4x }{ y }-y = 0 \rightarrow -4x(y)^{-1}-y=0 \rightarrow 4xy ^{-2} \frac{ dy }{ dx }-4y ^{-1}-\frac{ dy }{ dx }\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx }(\frac{ 4x }{ y ^{2} }-1) = -\frac{ 4 }{ y } \rightarrow \frac{ \frac{ 4 }{ y } }{ \frac{ 4x }{ y ^{2}-1 } } = \frac{ dy }{ dx } = \frac{ (y ^{2}-1) }{ xy }\]

OpenStudy (anonymous):

hopefully this is correct, it's been a while since I've done implicit differentiation.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

nope unfortunately there is nothing like this int he answers

OpenStudy (anonymous):

maybe instead of using the quotient rule, it would be easiest to start with \[y^2=-4x\] so you get \[2yy'=-4\] or \[y'=-\frac{2}{y}\]

OpenStudy (anonymous):

actually this is the second derivative started with 4x^2+y^2=9

OpenStudy (anonymous):

because the results here are not even close to what I have on the sheet

OpenStudy (anonymous):

found it!!! -(4x^2+16y^2)/y^3

OpenStudy (anonymous):

thanks anyway